# Bremsstrahlung Gamma Ray as a Probe of Short-Range Correlations in Nuclei

### v2

@@ Abstract
Atomic nuclei and dense nucleonic matter in neutron stars exhibit short-range correlations (SRCs), where nucleons form temporally correlated pair in close proximity. The impact of SRC on heavy-ion collisions (HICs) --which create dense nuclear matter in terrestrial laboratory-- has not been explored until now. In this letter, we present the neutron-proton bremsstrahlung $\gamma$-ray emission from HICs as a novel indicator of SRC in nuclei. By analyzing the hardening of the bremsstrahlung $\gamma$-ray spectrum, which results from the enhancement of high-momentum components above the Fermi surface in nucleon momentum distributions, we precisely determine the SRC fraction in nuclei to be $(20\pm 7)\%$ at $3\sigma$ confidence level. This experiment provides the first direct and accurate signature of SRC in low-energy HICs, offering a new method to study the parton distribution function of nucleons in nuclei.
Introduction -- Nucleonic matter, composed of strongly interacting protons and neutrons, exhibits complex many-body dynamics. Among these, some nucleons form short-range correlations (SRCs) [1, 2], which are temporal fluctuation effects that go beyond the mean-field approximation. SRCs have been experimentally observed and provide an explanation for the high-momentum tail (HMT) observed beyond the Fermi surface in nucleon momentum distributions [1, 3]. These high-momentum components significantly influence the properties of nuclear matter, particularly the density dependence of the nuclear symmetry energy [4, 5, 6], a quantity that is crucial in dense environments like neutron stars and heavy-ion collisions (HICs).
Studies indicate that SRCs are primarily characterized by neutron-proton ($np$) pairs in the isosinglet (T=0) channel at short distances, favoring $ud$ diquark isosinglet and spin-singlet configurations [7]. These structural features of SRCs have been shown to significantly influence the EMC effect [8] in nuclei [9, 10, 11], establishing a direct link between short-range nucleon correlations and quark-level modifications of nuclear structure [2, 12, 13, 14, 15, 16]. Furthermore, a recent global analysis based on the QCD factorization scheme has extracted the universal effective parton distribution within SRC pairs and the mean-field fraction using a variety of high-energy data [17]. These findings suggest that SRC studies may offer valuable insights into the modification of parton distribution functions (PDFs) in the nuclear environment.
@@
This letter reports on a new measurement of bremsstrahlung $\gamma$-rays from $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  collisions at 25 MeV/u with significantly improved statistics and extended energy range. The $\gamma$-ray spectrum is analyzed using the IBUU-MDI transport model, which incorporates the $np\rightarrow np\gamma$ process. The SRC-induced HMT is included in the initialization of the nucleon momentum. A precise $R_{\text{HMT}}$ is obtained, providing a new paradigm to study SRC of nucleons and their underlying partonic structure in nuclei through HICs.
Experiment-- The experiment was performed on Radioactive Ion Beam Line 1 (RIBLL1) at the Heavy Ion Research Facility in Lanzhou (HIRFL). The products from the reaction of 25 MeV/u $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  were measured by the Compact Spectrometer for Heavy Ion Experiment (CSHINE) [37, 38, 39]. Fig. 1 presents the detector setup. The charged isotopes of $Z\leq 5$ were measured by 6 silicon-strip detector telescopes (SSDTs) covering the polar angle $20^{\circ}<\theta_{\rm lab}<60^{\circ}$ in partial azimuth [40]. Neutrons are detected by a $4\times 5$ array of plastic scintillator installed at 2.0 meters to the target. With 4 fast $\rm BaF_{2}$ $\gamma$-detectors surrounding the target delivering starting time (T0) information, neutrons can be discriminated from the $\gamma$-rays [37, 41]. Besides, a liquid scintillator (LS) neutron detector in size of $\rm\phi 348~{}mm\times 1020~{}mm$ is mounted at 3 m to the target (not shown in Fig. 1).
@@
To measure the energetic bremsstrahlung $\gamma$-rays from the reactions on the target, a $4\times 4$ CsI(Tl)-based hodoscope has been mounted [42]. Each unit has a dimension of $70\times 70\times 250~{}\rm mm^{3}$. Three plastic scintillators were installed surrounding the CsI(Tl) hodoscope to veto the cosmic ray background as well as the $\gamma$ events with energy leaking to the outside. A dual-range amplifier was applied to record the small and large energy signals, and then the response of each CsI(Tl) unit was calibrated by a dual-range method [43]. The energy range of each individual CsI(Tl) unit was optimized in order to reach the maximum $\gamma$-ray energy deposit of 100 MeV in the hodoscope. To subtract the background, which mainly includes cosmic-ray muons and residual $\gamma$ radiation at low energies, the background spectrum was measured when the beam was off.
@@
Figure 4: (a) Comparison of the rebinned experimental $\gamma$ spectrum (black dots) in CM reference with various detector-filtered theoretical curves. Statistical (systematic) uncertainties are denoted by vertical bars (gray bands). (b) Likelihood distribution as a function of $R_{\rm HMT}$ with quadratic fit (dashed curve). The error bars represent the standard deviation of likelihood values by varying the calibration and detector response data sets. Confidence levels of $3\sigma$ and $5\sigma$ are indicated, respectively.
@@
The result reveals a new paradigm to probe the PDF of nucleon in nuclei (nPDF) via low energy HICs. According to the collinear factorization scheme [54], the nPDF parametrization is composed of a linear combination of PDFs of free nucleons and of SRC pairs, with the fraction fully encapsulated in the ratio of neutrons (protons) in SRC pairs $C_{n}^{\rm A}$ ($C_{p}^{\rm A}$). With $R_{\text{HMT}}=20\%$ in 124Sn ($A=124$, $\delta=0.194$), one writes $C_{n}^{124}=16\%$ and $C_{p}^{124}=24\%$, respectively, in accordance with the results obtained from high energy data [54]. The result also underscores the need for further systematic experimental studies of higher-energy $\gamma$-ray production in HICs. Interestingly, based on very similar mechanism of the presence of HMT, sub-threshold particle production in proton-nucleus collisions is proposed recently as a probe of SRC in nuclei [55].
Summary -- The precise measurement of the $np$ bremsstrahlung $\gamma$-rays in $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  at 25 MeV/u not only resolves a longstanding puzzle that emerged in the 1980s regarding the origin of high-energy $\gamma$-rays in HICs [56], but also introduces a new indicator of SRC, a significant phenomenon that is crucial to understand nuclear structure and the properties of dense nucleonic matter. By exploiting the advantage of HICs, where the $np$ bremsstrahlung $\gamma$-ray energy spectrum is hardened due to the HMT of nucleons resulting from SRC, we compare the experimental $\gamma$ spectrum with IBUU-MDI model predictions. This comparison allows us to precisely determine the high-momentum nucleon fraction in the 124Sn nucleus, $R_{\rm HMT}=(20\pm 7)\%$ at $3\sigma$ confidence level. Our experiment provides the first direct signature of SRC in low-energy HICs, establishing a new approach to study the parton momentum distribution of nucleons in nuclei.

### v3

# Precise Measurement of Short-Range Correlations in Nuclei from Bremsstrahlung Gamma Ray Emission in Low-Energy Heavy-Ion Collisions

@@ Abstract
Atomic nuclei and dense nucleonic matter in neutron stars exhibit short-range correlations (SRCs), where nucleons form temporally correlated pairs in proximity beyond mean-field approximation. It is essential to make precision measurement of the fraction of SRC since it carries the signature of underlying quark dynamics in nuclear medium. In this letter, we present the first high-precision measurement of neutron-proton bremsstrahlung $\gamma$-ray emission from the symmetric $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  reactions at 25 MeV/u. From the observed spectral hardening, the precise SRC fraction in the $\rm{}^{124}Sn$ nucleus is extracted to be $(20\pm 3)\%$. This result provides a novel, direct and unambiguous evidence of SRCs, and demonstrates that low-energy heavy-ion collisions offers a new approach to studying nuclear structure in connection with quark-level dynamics.
Introduction -- Nucleonic matter, composed of strongly interacting protons and neutrons, exhibits complex many-body dynamics. Among these, some nucleons form short-range correlations (SRCs) [1, 2], which are temporal fluctuation effects that go beyond the mean-field approximation. SRCs have been experimentally observed and provide an explanation for the high-momentum tail (HMT) observed beyond the Fermi surface in nucleon momentum distributions [1, 3]. These high-momentum components significantly influence the properties of nuclear matter, particularly the density dependence of the nuclear symmetry energy [4, 5, 6], a quantity that is crucial in dense environments like neutron stars and heavy-ion collisions (HICs). SRCs reflect the short-range structure of the nuclear force, and their precise measurement is key to uncovering the origin of nucleon-nucleon interactions.
Studies indicate that SRCs are primarily characterized by neutron-proton ($np$) pairs in the isosinglet (T=0) channel at short distances, favoring $ud$ diquark isosinglet and spin-singlet configurations [7]. These structural features of SRCs have been shown to significantly influence the EMC effect [8] in nuclei [9, 10, 11], establishing a direct link between short-range nucleon correlations and quark-level modifications of nuclear structure [2, 12, 13, 14, 15, 16]. Furthermore, a recent global analysis based on the QCD factorization scheme has extracted the universal effective parton distribution within SRC pairs and the mean-field fraction using a variety of high-energy data [17]. These findings suggest that SRC studies may offer valuable insights into the modification of quark dynamics in the nuclear environment, serving as a potential bridge between nuclear structure and the underlying partonic degrees of freedom.
@@
This letter reports on a new and high precision measurement of bremsstrahlung $\gamma$-rays from $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  collisions at 25 MeV/u with significantly enhanced statistics and extended high energy end of the spectrum. The $\gamma$-ray spectrum is analyzed using the IBUU-MDI transport model, which incorporates the $np\rightarrow np\gamma$ process. The use of a symmetric reaction allows for an unambiguous extraction of $R_{\text{HMT}}$ in the $\rm{}^{124}Sn$ nucleus.
Experiment-- The experiment was performed on Radioactive Ion Beam Line 1 (RIBLL1) at the Heavy Ion Research Facility in Lanzhou (HIRFL). The products from the reaction of 25 MeV/u $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  were measured by the Compact Spectrometer for Heavy Ion Experiment (CSHINE) [37, 38, 39]. Fig. 1 presents the detector setup. The charged isotopes of $Z\leq 5$ were identified and reconstructed by 6 silicon-strip detector telescopes (SSDTs) covering $20^{\circ}<\theta_{\rm lab}<60^{\circ}$ in partial azimuth [40, NuclSciTech.36.132]. Neutrons are detected by a $4\times 5$ array of plastic scintillator installed at 2.0 meters to the target. With 4 fast $\rm BaF_{2}$ $\gamma$-detectors surrounding the target delivering starting time (T0) information, neutrons can be discriminated from the $\gamma$-rays [37, 41]. Besides, a liquid scintillator (LS) neutron detector of $\rm\phi 348~{}mm\times 1020~{}mm$ is mounted at 513 cm to the target (not shown in Fig. 1).
@@
To measure the energetic bremsstrahlung $\gamma$-rays from the reactions on the target, a $4\times 4$ CsI(Tl)-based hodoscope has been mounted [42]. Each unit has a dimension of $70\times 70\times 250~{}\rm mm^{3}$. Three plastic scintillators were installed surrounding the CsI(Tl) hodoscope to veto the cosmic ray background as well as the $\gamma$ events with energy leaking to the outside. A dual-range amplifier was applied to record the small and large energy signals, and then the response of each CsI(Tl) unit was calibrated by a dual-range method [43]. The energy range of each individual CsI(Tl) unit was optimized in order to reach the maximum $\gamma$-ray energy deposit of 100 MeV in the hodoscope. The background, which mainly includes cosmic-ray muons and residual $\gamma$ radiation at low energies, was measured when the beam was off.
@@
Figure 4: (a) Comparison of the rebinned experimental $\gamma$ spectrum (black dots) in CM reference with various detector-filtered theoretical curves. Statistical (systematic) uncertainties are denoted by vertical bars (gray bands). (b) Likelihood distribution as a function of $R_{\rm HMT}$ with quadratic fit (dashed curve). The error bars represent the standard deviation of likelihood values by varying the calibration and detector response data sets. Several main confidence levels are indicated.
@@
The result reveals a potential new paradigm to probe the PDF of nucleon in nuclei (nPDF) via low energy HICs. According to the collinear factorization scheme [54], the nPDF parametrization is composed of a linear combination of PDFs of free nucleons and of SRC pairs, with the fraction fully encapsulated in the ratio of neutrons (protons) in SRC pairs $C_{n}^{\rm A}$ ($C_{p}^{\rm A}$). With $R_{\text{HMT}}=20\%$ in 124Sn ($A=124$, $\delta=0.194$), one writes $C_{n}^{124}=16\%$ and $C_{p}^{124}=24\%$, respectively, in accordance with the results obtained from high energy data [54]. The result also underscores the need for further systematic experimental studies of higher-energy $\gamma$-ray production in HICs. Interestingly, based on very similar mechanism of the presence of HMT, sub-threshold particle production in proton-nucleus collisions is proposed recently as a probe of SRC in nuclei [55].
Summary -- This work presents the first direct and quantitative evidence of short-range correlations (SRCs) in low-energy heavy-ion collisions, based on the highest-quality measurement to date of neutron-proton bremsstrahlung $\gamma$-rays up to 100 MeV in symmetric $\rm{}^{124}Sn$+$\rm{}^{124}Sn$  reactions at 25 MeV/u. This resolves a longstanding puzzle from the 1980s regarding the origin of high-energy $\gamma$-rays in such collisions [56], and establishes bremsstrahlung emission as a novel and sensitive probe of SRCs in nuclei. By analyzing the spectral hardening caused by high-momentum nucleons, we determine the SRC fraction in $\rm{}^{124}Sn$ to be $R_{\rm HMT}=(20\pm 3)\%$, through comparison with transport model predictions. This result surpasses the previous experimental attempts in both precision and clarity, made possible by improved statistics, extended spectral coverage, and a symmetric reaction system. Our findings open a new avenue for studying nucleon SRCs and their possible connection to partonic dynamics in nuclei through low-energy heavy-ion collisions.